OpenStudy (fairy733189):

A basketball is thrown upwards. The height f(t), in feet, of the basketball at time t, in seconds, is given by the following function: f(t) = -16t^2 + 16t + 32 Which of the following is a reasonable domain of the graph of the function when the basketball falls from its maximum height to the ground? 0.5 < t < 1 0.5 < t < 2 1 < t < 2 1 < t < 1.5

OpenStudy (fairy733189):

@math&ing001

OpenStudy (math&ing001):

Same, you find the vertex

OpenStudy (fairy733189):

the vertex is 1/2 but I dont see that as an option

OpenStudy (math&ing001):

Then you find the time when the basketball hits the groung (aka solve f(t)=0)

OpenStudy (math&ing001):

1/2 = 0.5

OpenStudy (fairy733189):

Oh ok let me try that

OpenStudy (fairy733189):

sorry I did't see the 0.5

OpenStudy (fairy733189):

f(0) = -16(0)^2 + 16(0) + 32 f(0) = 32

OpenStudy (math&ing001):

no f(t)=0 and you solve for t

OpenStudy (fairy733189):

what do you mean. Can you explain that please?

OpenStudy (math&ing001):

f(t)=0 ie -16t^2 + 16t + 32 = 0 We solve the quadratic equation

OpenStudy (fairy733189):

oh ok I learned this in my class but it was kid of confusing so can you explain how I am supposed to solve this?

OpenStudy (fairy733189):

Im sorry im really not that great at math

OpenStudy (math&ing001):

You might wanna check this http://www.purplemath.com/modules/quadform.htm

OpenStudy (fairy733189):

so would I substitute the numbers into the quadratic formula?

OpenStudy (math&ing001):

Yep !

OpenStudy (fairy733189):

ok let me try: x = -16 ± √16^2 - 4(-16)(32) / 2(-16) x = -16 ± √256 - (-2048) / -32 x = -16 ± √2304 / -32 x = -16 ± 48 / -32 x = -16 + 48 / -32 And -16 - 48 / -32 x = -1 AND x = 2

OpenStudy (math&ing001):

Correct but time can't be negative so x=2

OpenStudy (fairy733189):

ok so the answer would be option 2?

OpenStudy (math&ing001):

Correct

OpenStudy (fairy733189):

Thank You so much!!!

OpenStudy (math&ing001):

Welcome !!